1. Find the volume of the solid under the cone z= sqrt (x^2 + y^2) and over the ring 4 |\eq x^2 + y^2 |\eq 25.

2. Find the volume of the solid under the plane 6x + 4y + z= 12 and over the disk with border x^2 + y^2 = y.

3. The area of the smallest region, locked by the spiral r\Theta= 1, the circles r=1 and r=3 and the polar axis.
Answer #1

Colin is 2 Given = Tarty and ginen dist x² + y² s 4 So find the volume under the Cone abowe the disk, the polar coordinates.the Sol- to find volume by double integral founded by plane: Gx+4y +z = 12 l.e., 12-62- Ay . R = of 4 5 JJ Z dA IS (12- 6x-4y3 Yol -2 Ginen 10 = 1, circles 91=1,96=3 ISS 3 and 91 isos š Aoua og plana - (oko *.103 • 3-1 2 with double The other way I